OpenStudy (anonymous):

The manufacturing cost of a cell phone was $124. Sellers first marked up the cost by 20 percent. Sales were poor, so sellers decreased the markup to 15 percent. Calculate the old and new selling prices.

OpenStudy (anonymous):

124 marked up 20% : 124 + 124(.20) = 124 + 24.80 = 148.80 ==old selling price Now they decreased the mark down by 15% : 148.80 - 148.80(.15) = 148.80 - 22.32 = 126.48 ==new selling price

OpenStudy (anonymous):

OMG thank you i was stuck on this problem for a while

OpenStudy (anonymous):

glad to help :)